OpenStudy (anonymous):

Eliminate the parameter. x = 4 cos t, y = 4 sin t ?

OpenStudy (anonymous):

would it be x^2 = 4 cos^2 (t) y^2 = 4sin^2 (t) ??

OpenStudy (psymon):

You're on the right track, but x^2 = 16cos^2(t) and y^2 = 16sin^2(t)

OpenStudy (anonymous):

oh i forgot to square the 4. Is there any step after that? @Psymon

OpenStudy (psymon):

Well, the parameter is still there, all you did was manipulate it. Try adding them and see what you can make happen :3

OpenStudy (anonymous):

16cos^2(t)+16sin^2(t) ? @psymon

OpenStudy (psymon):

Mhm. Now you got an identity going on there.

OpenStudy (anonymous):

@Psymon is that the final answer?

OpenStudy (psymon):

Well, if I factor out 16, I have 16(sin^2(x) + cos^2(x)). Do you recognize that identity inside of the parenthesis?

OpenStudy (psymon):

In the end, you're trying to do two things. One, you want one equation, not 2. Secondly, you want to eliminate t. We've manipulated the problem, but t is still there. If you use an identity on that sin and cos, though, t will disappear, given you recognize the identity.

OpenStudy (anonymous):

the identity is equal to 1? @Psymon

OpenStudy (psymon):

Yep. So in the end, you have x^2 + y^2 = 16, and that would be your answer.

OpenStudy (anonymous):

@Psymon thank you!